Sol Summary Report – Feb 19th

CREW 205 Sol Summary

Sol: 9

Sol Summary Title: A Martian Day in the Hab

Author: Ghanim Alotaibi

Mission status: Nominal

Sol activity summary: The bad weather last night affected today’s activities. EVA plans were cancelled last night, and astronomical observations submitted a few days ago needed to wait for an extra night. This morning the crew woke up with an exceptional white view of the Martian landscape from the Hab window. While having breakfast, the white bright snow was melting down, and as we approached lunch time, the muddy Martian land started to finally appear.

Duties in the Hab, and its facilities started just after finishing breakfast as planned last night. Dave, Nathan and Veronica went to the Science Dome to clean, measure and characterize the samples. In fact, those samples made the whole crew so excited, and was the main topic of today and last night. Our XO today, Hannah, prepared lunch for the crew with the help of Dan. Maria, continued to work hard for our outreach movie, this time, with the help of Natalia. Natalia started preparing a guideline document for future crews and discussed with the commander of the day, Ghanim.

In addition to the above work, the crew continued the normal daily maintenance, cleaning of the Hab and writing reports as usual. The only new duty was to sweep out the snow accumulated in the tunnel to avoid muddy walk way.

In the night, Hannah cooked us a delicious fried rice with chicken, and the crew prepared to watch the Apollo 13 Film.

Anomalies: EVA cancelled

Weather: partly cloudy

Reports to be filed: Sol Summary, Operations Report, Greenhab Report, Journalist Report.

Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site